OpenStudy (anonymous):

Evaluate the given interval by using the given substitution. (dx/((8x+4)^4)) where u=8x+4

OpenStudy (anonymous):

\[u=8x+4\]\[du=8dx\]\[\frac{1}{8} \int\limits_{?}^{?} du/u\]

OpenStudy (anonymous):

\[= \frac{1}{8}\int\limits{du/u ^{4}}\]

OpenStudy (anonymous):

how do i write that out? is it 1/8(8/(8x+4)^4)?